# Post-detection inference for sequential changepoint localization

### v2

@@ A.3 Omitted proofs from Section 4
    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\},\tau<T)\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau<T),$
@@
    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L}^{\alpha r_{t}})$
    $=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)$
@@
  2. (b)
    For $\mathcal{C}_{2}$: Proof is similar to part (a). Here we have for fixed $r_{t}=1$ with $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverages $1-\beta$ and $1-\gamma$ respectively,
    $P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta+\gamma.$
    Similarly, we have
    $P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
    If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\geq 1-\alpha-\beta$, we obtain
    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta+\gamma}{\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)}\geq 1-\frac{\alpha+\beta+\gamma}{\mathbb{P}_{\theta_{0},\infty}(\tau=\infty)}\geq 1-\frac{\alpha+\beta+\gamma}{1-\delta}.$
    Similarly,
    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)-(\alpha+\beta+\gamma)\geq 1-\alpha-\beta-\gamma-\delta.$
Appendix B Another concrete example with Poisson distribution
B.1 Known pre-change and composite post-change
@@
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (2.8), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let $\tau^{\theta}_{j}$ be the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark B.1) and they are finite. So, we can choose $L=\infty$.

### v3

@@ Appendix A Omitted proofs
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}^{*}}\},\tau<T)\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau<T),$
@@
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L}^{\alpha r_{t}^{*}})$
$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}^{*}}\})$
$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}^{*}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)$
@@
Proof of Theorem 5.1.
First note that $\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CS}(X_{T},\cdots,X_{\tau};1-\eta_{1}r_{T}^{*})\mid r_{T}^{*}\right)\leq\eta_{1}r_{T}^{*}$ and $\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{0}\notin\displaystyle\operatorname{CI}(X_{1},\cdots,X_{T-1};1-\eta_{0}r_{T}^{*})\mid r_{T}^{*}\right)\leq\eta_{0}r_{T}^{*}$, because $r_{T}^{*}$ is independent of $X_{i}$’s. Now,
$\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})\mid\tau\geq T\right)$
$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*}),T\in\mathcal{C}\mid\tau\geq T\right)+\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(T\notin\mathcal{C}\mid\tau\geq T\right)$
$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CS}(X_{T},\cdots,X_{\tau};1-\eta_{1}r_{T}^{*})\mid\tau\geq T\right)+\alpha$
$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CS}(X_{T},\cdots,X_{\tau};1-\eta_{1}r_{T}^{*})\right)/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+\alpha$
$=\mathbb{E}(\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CS}(X_{T},\cdots,X_{\tau};1-\eta_{1}r_{T}^{*})\mid r_{T}^{*}\right))/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+\alpha$
$\leq\mathbb{E}(\eta_{1}r_{T}^{*})/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+\alpha$
$=\alpha+\eta_{1}\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq T)/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$
$\leq\alpha+\eta_{1},$
where the last inequality follows from Assumption 3.3. Similarly,
$\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})\mid\tau\geq T\right)$
$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*}),T\in\mathcal{C}\mid\tau\geq T\right)+\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(T\notin\mathcal{C}\mid\tau\geq T\right)$
$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CI}(X_{1},\cdots,X_{T-1};1-\eta_{0}r_{T}^{*})\mid\tau\geq T\right)+\alpha$
$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CI}(X_{1},\cdots,X_{T-1};1-\eta_{0}r_{T}^{*})\right)/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+\alpha$
$=\mathbb{E}(\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CI}(X_{1},\cdots,X_{T-1};1-\eta_{0}r_{T}^{*})\mid r_{T}^{*}\right))/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+\alpha$
$\leq\mathbb{E}(\eta_{0}r_{T})/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+\alpha$
$\leq\alpha+\eta_{0}.$
Appendix B Computationally and statistically efficient subroutines in Algorithm 2
$L_{t}(\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (B.1)
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})}.$ (B.2)
$M_{t}^{(n)}(X_{1},\cdots,X_{n})=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (B.3)
Note that the test statistic $M_{t}$ for testing $\mathcal{H}_{0,t}$ reduces to
$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{if }t\leq\tau<\infty,\\ -\infty,&\text{if }t>\tau,\text{ or }\tau=\infty.\\ \end{cases}$ (B.4)
$Y_{n}^{t}(x,\theta):=\begin{cases}G(x,\theta_{0}),&\text{if }n<t\\ G(x,\theta),&\text{ otherwise. }\end{cases}$ (B.5)
Let us provide a few examples of the function $G$. If the distributions in the parametric family are continuous, the inverse distribution function $F_{\theta}^{-1}$ exists and it is easy to verify that $G(x,\theta)=F_{\theta}^{-1}(F_{\theta^{\prime}}(x))$ works. In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$. In the case of discrete distributions, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta)\}_{n}$ across all $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution.
Let us now fix $t\in\{1,\cdots,\tau\}$. Computing the detection time $\tau_{j,t}^{\theta}$ for each $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether $\text{there exists }\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}^{\prime}_{\tau-t+1}$ is a finite set. However, it is often straightforward to compute $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$. Also, we can often derive a suitable upper bound on the threshold, which is computationally and statistically efficient. For example, we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$. Again, we bound $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)$ by $V_{t,L}^{j}$, where for $L<\infty$,
$V_{t,L}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j},\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{if }t\leq t_{2}^{j}\leq L,\\ \infty,&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}.$ (B.6)
And for $L=\infty$,
$V_{t,\infty}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty,\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{otherwise,}\\ \end{cases}$ (B.7)
which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, see Section B.1). We define the confidence set as
$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (B.8)
We observe that $\mathcal{C}_{1}^{\prime}\supseteq\mathcal{C}_{1}$, where $\mathcal{C}$ is the set produced by Algorithm 2, since
$\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{V_{t,L}^{j}\}_{j=1}^{B})$ $\geq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$
$\geq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}).$
Corollary B.1.
$\mathcal{C}^{\prime}_{1}$ defined in (B.8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.4.
Next, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with the Poisson distribution in Section G.3.
B.1 A concrete example: Gaussian mean shift problem
$\mathcal{S}^{\prime}_{\tau-t+1}=\left[\bar{X}_{t:\tau}-\frac{s_{\tau-t+1}(\beta)}{\sqrt{\tau-t+1}},\bar{X}_{t:\tau}+\frac{s_{\tau-t+1}(\beta)}{\sqrt{{\tau-t+1}}}\right]\cap\Theta_{1},$ (B.9)
$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\epsilon_{n}^{j}+\theta_{0},\text{ if }n<t,\\ \epsilon_{n}^{j}+\theta,\text{ otherwise, }\end{cases}$ (B.10)
where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$. We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (B.4), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let, $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark B.2) and they are finite. So, we can have $L=\infty$.
$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{(t-i)(\theta-\theta_{0})}{t^{\prime}-i+1}\right)^{2}+\text{const},\text{ if }i<t,\\ \text{const},\text{ if }i=t,\\ \displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}+\theta-\theta_{0}\right)^{2}+\text{ const},\text{ if }i>t,\end{cases}$
Remark B.2.
[…]
$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (B.11)
[…]
For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$. Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{n}(2X_{i}-\theta_{0}-\theta)\}$ increases as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Similarly, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Also, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\neq\theta_{0}\}$, if we employ Gaussian weights (it is conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution. One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section G.2.
The Gaussian mean shift example we provided serves as a concrete and illustrative case to clarify the methodology. It is worth noting that $C^{\prime}$, as defined in (B.8), is practically implementable as long as the following relatively weak conditions are met:
  1. 1.
    The coupling function $G$ is available (which is always the case when distributions are continuous and in the case of discrete distributions, even if a suitable $G$ is not directly available, our method may still be implementable, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution).
  2. 2.
  3. 3.
  4. 4.
    If we employ the likelihood-ratio-based test statistic (B.4), then we must be able to maximize $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
B.2 Another concrete example: Poisson rate change problem
@@
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in Section 3, with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let $\tau^{\theta}_{j}$ be the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark B.3) and they are finite. So, we can choose $L=\infty$.